Combustion of 0.2g carbon-hydrogen-oxygen compound yields 0.2998g CO2 and 0.0819g H2O. Find percentage composition and empirical formula of the sample
1
Expert's answer
2020-12-10T07:44:57-0500

Let's detemine masses of carbon, hydrogen and oxygen in 0.2 g of CxHyOz

m(C) = 0.2998*12/44 = 0.0818 g

m(H) = 0.0819*2/18 = 0.00910 g

m(O) = 0.2-0.0818-0.0091 = 0.1091 g

Let's detemine mass fractions of carbon, hydrogen and oxygen in CxHyOz

"\omega"(C) = 0.0818/0.2 = 0.409 = 40.9%

"\omega"(H) = 0.00910/0.2 = 0.0455 = 4.55%

"\omega"(O) = 0.1091/0.2 = 0.5455 = 54.55%

Let's find empirical formula of CxHyOz:

x : y : z = 0.0818/12 : 0.00910 : 0.1091/16 = 0.0068 : 0.0091 : 0.0068 = 1 : 1.33 : 1 = 3 : 4 : 3

C3H4O3